The length of an elastic string is X m when the tension is 8N and Y m when tension is 10 N. The length in metres when the tension is 18N isi) 4X-5Yii)5Y-4Xiii)9X-4Yiv)4Y-9X

Grade:11

1 Answers

jitender
6 years ago
T=kxNow ◇T=k◇x10-8=k (y-x) ...1st18-8=k (z-x) .....2nd2nd÷1st10/2=(z-x)/(y-x)5y-5x=z-xZ=5y-4xSo length of string is 5y-4x when T=18hence ii option is correct